OpenStudy (anonymous):

how do you make y a function of x when 4.2x-2y=16.8? would it be -2y=16.8 -4.2x

OpenStudy (anonymous):

nearly divide thru by -2: so y = -8.4 + 2.1x

OpenStudy (anonymous):

ohh ok so you compleatly isolate y.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so if 8x-5x+21=36-6y, its y=2x-15? sorry this is the last q